The analysis of an unknown compound shows that it is made up of 14.14% C, 2.37% and 83.48% O. What is the chemical formula of the compound if it’s molar mass is 84.93g/mol?


1
Expert's answer
2021-06-07T03:24:06-0400

Solution:

Most likely there is a mistake in the task. And 83.48% corresponds to the percentage of chlorine (Cl).


Molar mass of carbon (C) is 12.0107 g mol-1

Molar mass of hydrogen (H) is 1.00784 g mol-1

Molar mass of chlorine (Cl) is 35.453 g mol-1


Assume that the sample has a mass of 100 g.

Convert %values to grams:

Mass of C = w(C) × Mass of sample = 0.1414 × 100 g = 14.14 g

Mass of H = w(H) × Mass of sample = 0.0237 × 100 g = 2.37 g

Mass of Cl = w(Cl) × Mass of sample = 0.8348 × 100 g = 83.48 g


Convert to moles:

14.14 g C × (1 mol C / 12.0107 g C) = 1.1773 mol C

2.37 g H × (1 mol H / 1.00784 g H) = 2.3516 mol H

83.48 g Cl × (1 mol Cl / 35.453 g Cl) = 2.3547 mol Cl


Divide all moles by the smallest of the results:

C: 1.1773 / 1.1773 = 1.000 = 1

H: 2.3516 / 1.1773 = 1.997 = 2

Cl: 5.2178 / 1.1773 = 2.000 = 2


Therefore, the empirical formula of an unknown compound is CH2Cl2

M(CH2Cl2) = Ar(C) + 2×Ar(H) + 2×Ar(Cl) = 12.0107 + 2×1.00784 + 2×35.453 = 84.9324 (g/mol)

Therefore, the chemical formula of the compound corresponds to the empirical formula.

The chemical formula of the compound is CH2Cl2


Answer: The chemical formula of the compound is CH2Cl2